ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30, 47 Bredimas, A. P., University Paris 7, France, and NCSR Demokritos Attiki-Greece Schlicht functions and exotic sprectral analysis: Applications to some conjectures and problems in analysis Recallings on “exotic stropharmonic analysis” and “exotic spectral analysis” are given. Applications to the class of schlicht functions are considered: consequences in Dirichlet series and and in the area of the Bieberbach-Rogosinski-Sheil-Small conjecture are reported. Reference: [1] Selected Conference Texts . .., ISBN 960 8492-06-8 Athens’s Centre, Greece, 1994, 205p; [2] Preprints 1985 - 1996. .., ISBN 960 8492-38-6 Athens’s Centre, Greece, 1997, 246p: [3] Generalized Analytical Functional. .., 9p. Preprint 1995; [4] Analyse Spectrale Exotique. .. , 6p. Preprint 1997.